daniel wants to make a 2% saline solution. He has already poured 12ml of pure water into a beaker. how many ml of a 8% saline solution must he add to this to create the desired mixture?

Found 2 solutions by josgarithmetic, stanbon:
Answer by josgarithmetic(39617)
Answer by stanbon(75887)

daniel wants to make a 2% saline solution. He has already poured 12ml of pure water into a beaker. how many ml of a 8% saline solution must he add to this to create the desired mixture?
------
Equation:
water + water = water
12 + 0.92x = 0.98(12+x)
-------------
12 + 0.92x = 0.98*12 + 0.98x
-----
0.02*12 = 0.06x
x = 4 ml
-------------
Cheers,
Stan H.
--------------